# Gesplat: Robust Pose-Free 3D Reconstruction via Geometry-Guided Gaussian Splatting

### v1

@@ Abstract
Figure 1: Novel View Synthesis Comparisons.We introduce Gesplat, an efficient framework for novel view synthesis and 3D reconstruction from sparse-view unposed inputs.Compared with other pose-free methods, our method achieves higher PSNR and reconstructs more consistent geometry with finer details.
I Introduction
@@
Neural Radiance Fields (NeRF) [1], an implicit neural representation that has shown remarkable success in reconstructing scenes and synthesizing photorealistic novel views. To improve performance under sparse views, NeRF-based methods have attempted in the aspect of pretraining [2, 3, 4], regularization terms [5, 6], and external priors [7, 8, 9, 10], achieving significantly improved results. Nonetheless, these approaches still suffer from high training and rendering costs.
@@
In sparse-view settings, limited input data results in inconsistent scene geometry, and the high interdependence among Gaussian attributes forces the optimization to make a trade-off between optimizing shape and position. Some optimization-based methods [22, 23, 24] excel in real-time scene reconstruction, but when applied to sparse-view scenarios, they usually introduce invalid geometric priors, leading to incorrect topology and scale ambiguity. Incorporating additional geometric priors is essential to constrain scene structure and improve the accuracy of 3D point positions. Depth information is a commonly used prior due to its connection between 2D images and 3D structure. Some researches [25, 26] incorporate monocular depth priors [27] as regularization, yet such handcrafted constraints often suffer from scale inaccuracy and multi-view inconsistency, which can adversely affect final rendering quality.
In this paper, we propose Gesplat, an effective framework that incorporates appropriate geometric priors to constrain the scene structure while introducing optimization and regularization techniques to refine the scene details. Inspired by [28], we adopt a hybrid Gaussian representation combining ordinary and ray-based Gaussians leveraging matching priors. By binding Gaussians to matching rays, we enhance multi-view consistency and preserve geometric structure. Furthermore, we introduce dual optimization of position and shape to stabilize training and ensure accurate surface convergence. After obtaining a relatively accurate scene representation, we employ a Graph Neural Network (GNN) [29, 30] to further optimize Gaussian attributes. The Gaussians, treated as vertices, are connected based on spatial neighbor relationship, and learnable offsets are applied to update their attributes. To achieve more accurate depth estimation, we utilize an explicit depth computation method within an epipolar geometry framework through optical flow, following [31]. The estimated reliable depth maps are used sequentially for regularization, further improving the rendering quality. During training, we preserve joint training from [19] to optimize the Gaussian parameters and camera poses at the same time. Novel view synthesis comparisons with other pose-free methods are shown in Fig. 1, which exhibit the superior performance of our method.
@@
  • We introduce a hybrid Gaussian representation with dual optimization of position and shape based on matching priors.
  • We design a graph-guided optimization module to refine Gaussian attributes for detailed scene recovery.
  • We apply flow-based depth regularization to improve the quality of rendered images during training.
@@
II Related Work
II-A 3D Reconstruction
@@ II Related Work
As a fundamental challenge in computer vision, 3D reconstruction aims to recover the 3D geometric structure of a scene from a set of input 2D images. Traditional optimization-based methods, such as Structure from Motion (SfM) [15] and Multi-View Stereo (MVS) [32, 33, 34], rely on rigorous geometric principles to estimate accurate camera parameters and dense point clouds. COLMAP, which integrates SfM and MVS, is one of the most popular and powerful open source tools. However, its high computational cost, sensitivity to errors in SfM-estimated camera poses and sparse point clouds, and dependence on dense viewpoint coverage limit its applicability in textureless regions and areas with low overlap.The advent of deep learning has spurred the development of learning-based reconstruction methods to overcome some limitations of traditional pipelines. As a pioneering work, MVSNet [35] constructs a 3D cost volume and uses 3D convolutional neural networks to regress depth maps. Follow-up works such as R-MVSNet [36] alleviate memory consumption through recurrent networks, while CVP-MVSNet [37] introduces a coarse-to-fine pyramid structure to improve efficiency and accuracy. Nonetheless, these approaches typically assume that camera poses are known and accurate, which is usually unreliable in real-world scenarios. More recently, DUSt3R [20] notably proposes an end-to-end model for feed-forward 3D reconstruction. Unlike traditional methods that require camera calibration or viewpoint poses, DUSt3R leverages a Transformer architecture to directly estimate 3D point clouds from unconstrained image pairs, followed by global alignment to produce a complete scene representation. Its sibling, MASt3R [38], augments DUSt3R with a second network head to generate dense local features and introduces a novel matching loss for constraint. To avoid the computational cost of iterative post-optimization required by DUSt3R, VGGT [21] introduces a feed-forward network that directly generates more continuous point clouds and accurate camera poses, capable of processing more than two images simultaneously. VGGT achieves state-of-the-art reconstruction performance compared to DUSt3R and MASt3R. Inspired by this, we adopt VGGT for dense scene initialization.
II-B Novel View Synthesis
@@
Novel view synthesis (NVS) aims to render high-quality images from unseen viewpoints. Neural Radiance Fields (NeRF) [1], which implicitly represent scenes using multi-layer perceptions (MLPs) and synthesize images via volume rendering, has become one of the most influential NVS methods. Despite achieving high-quality reconstructions, NeRF suffers from high computational costs, leading to slow training and inference. Subsequent research has improved NeRF in various aspects, $e.g.$, quality [39, 40, 41, 42, 43], efficiency [44, 45, 46], and pose-free [47, 16, 48, 17]. More recently, 3D Gaussian Splatting (3DGS) [11] explicitly represents scenes using anisotropic 3D Gaussians [49], enabling rapid reconstruction and real-time rendering while maintaining high visual quality. 3DGS has been extensively explored in many aspects, such as [23, 50, 14, 51, 52, 53, 54, 55], demonstrating superior performance compared with NeRF-based approaches. However, conventional NVS methods heavily rely on dense input views and accurate camera poses derived from SfM, which is often impractical to attain, leading to significant performance degradation under sparse view settings.
II-C Sparse-View Novel View Synthesis
@@
Sparse-view novel view synthesis addresses a practical but more challenging task of synthesizing novel views from a limited set of input images, where overfitting occurs due to insufficient supervision. To mitigate these issues, various strategies have been proposed. Methods like [56, 57, 58] apply regularization techniques to reduce floating artifacts and enhance geometric consistency. [7, 9] incorporate depth priors from SfM or monocular depth estimation to supervise scene geometry, thereby reducing geometric errors and improving rendering image quality. Generalizable NeRF models like [4, 59] enable feed-forward inference, allowing fast adaptation to new scenes and improving practicality. Despite these advancements, NeRF-based methods remain computationally intensive due to volume rendering.In contrast to NeRF-based methods, several studies combined with 3DGS attempt to eliminate reconstruction inconsistency. To better capture spatial structure, [60, 26, 25, 61] introduce depth information into the explicit 3DGS representation. However, inaccuracies in monocular depth estimation often degrade reconstruction quality in sparse-view scenarios. [13, 62, 63] focus on robust scene initialization and removal of incorrectly positioned Gaussians. Some methods [63, 64, 65] introduce various regularization strategies to mitigate geometric degradation and suppress rendering artifacts. Alternative scene representations, such as [66, 28], are designed to produce accurate and detailed 3D scenes with enhanced geometry. To better preserve geometric structure completeness and detail, we introduce a hybrid Gaussian representation to learn scene representation, with flow-based depth regularization and graph-guided optimization incorporated to refine reconstruction results in weakly textured regions.
@@
Figure 2: Overall framework of Gesplat. Given a few input images, we first generate dense point clouds and camera poses from VGGT, extract matching priors and predict optical flow. Subsequently, we randomly initialize hybrid gaussian representation, using ray-based Gaussians to optimize the gaussian position. Graph-guided optimization and joint optimization are then applied to refine the Gaussians and camera poses. Finally, we employ flow-estimated depth and matching depth as rendering geometry regularization.
III Method
@@
III-A Preliminary
Gaussian Splatting. 3DGS[11] explicitly represents a scene through anisotropic 3D Gaussians, achieving high-quality scene reconstruction and real-time rendering. Each Gaussian is parameterized by: a position vector $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, spherical harmonic (SH) coefficients, and an opacity $\alpha\in[0,1)$. The influence for a point $x$ in 3D space is defined as:
@@ III-A Preliminary
Dense Initialization. Although 3DGS is efficient in real-time rendering, its initialization relies on a sparse point cloud and accurate camera parameters from Structure-from-Motion (SFM) [15]. This usually leads to inadequate information in sparse-view scenarios, which causes overfitting on training views and overly smooth textures. Thanks to advances in deep learning, recent learning-based 3D reconstruction methods bypass the need for explicit keypoint extraction and matching with traditional methods in an end-to-end way. Among these, DUSt3R [20] estimates 3D point clouds directly from 2D image pairs without precise camera calibration, and employs a global alignment strategy for multi-view 3D reconstruction. However, its iterative post-optimization steps results in high computation cost. To address this, we utilize VGGT [21], a feedforward neural network that produces high-fidelity dense point clouds and accurately estimated camera poses even from sparse and low-overlap datasets. This ensures geometric continuity in the reconstructed scene and provides a more robust initialization for 3DGS.
III-B Hybrid Gaussian Representation and Optimization
@@
Matching Priors. Matching priors refer to ray correspondence and ray position. 1) Ray Correspondence: A pair of matching rays $\{r_{i},r_{j}\}$ from view $i$ and view $j$ corresponds to the same 3D point. Given image $I_{i}$ and $I_{j}$, matching ray pair $\{r_{i},r_{j}\}$, pixel points $\{p_{i},p_{j}\}$, intrinsic parameters $\{K_{i},K_{j}\}$, and extrinsic parameters $\{[R_{i},t_{i}],[R_{j},t_{j}]\}$, we assume that each ray intersects the surface at point $X_{i}$ and $X_{j}$, satisfying $X_{i}=X_{j}$. Thus we can get the projection from view $i$ to view $j$ as:
@@
Ray-based Gaussian. Given an image pair $I_{i}$ and $I_{j}$ with $N$ matching ray pairs $\{r_{i}^{k},r_{j}^{k}\}_{k=1}^{N}$, there exist $N$ corresponding pairs of ray-based Gaussians $\{G_{i}^{k},G_{j}^{k}\}_{k=1}^{N}$. Each primitive has similar attributes with ordinary Gaussian but differs in position representation. The new representation of position $mu^{\prime}$ is given as follow:
@@
Position Optimization. Leveraging the binding strategy between ray-based Gaussians and their matching rays, we optimize Gaussian positions using the projection relation. For a matching ray pair $\{r_{i},r_{j}\}$ with corresponding Gaussians $\{G_{i},G_{j}\}$ at positions $\mu_{i}^{\prime}=o_{i}+z_{i}d_{i}$ and $\mu_{j}^{\prime}=o_{j}+z_{j}d_{j}$, the projection error is defined as:
@@
Rendering Geometry Optimization. In addition to position, other inaccurate attributes can also affect rendering quality. To address this, we optimize the rendering geometry. Through Eq. (4), we obtain the rendering depth maps $D_{i}$ and $D_{j}$. Then we use the depth values $\{D_{i}(p_{i}),D_{j}(p_{j})\}$ at matching pixels to yield corresponding 3D points:
@@
III-C Graph-guided Optimization
Considering the view inconsistency under a sparse-view setting, we introduce a Graph Neural Network (GNN) to better learn the spatial geometry structure, improving the scene reconstruction quality. We utilize a graph $\mathcal{G=(V,E)}$ to capture complex geometric features for Gaussians, where vertices $\mathcal{V}=\{v_{i}\}_{i=1}^{M}=\{z_{i},q_{i},s_{i},c_{i},\alpha_{i}\}_{i=1}^{M}$ represent the Gaussian attributes, and edges $\mathcal{E}$ describe the spatial adjacency based on K nearest neighbors (KNN). The K nearest neighbors of point $\mu_{i}$ are calculated as:
@@ III-C Graph-guided Optimization
where $r$ is a radius threshold.
Upon generating the graph network, we formulate its output as follows:
@@
$z_{i}^{\prime}$ $=z_{i}+\lambda_{z}\cdot\Delta z_{i}$ (14)
$s_{i}^{\prime}$ $=s_{i}+\lambda_{s}\cdot\Delta s_{i}$
$r_{i}^{\prime}$ $=s_{i}+\lambda_{r}\cdot\Delta r_{i}$
$c_{i}^{\prime}$ $=c_{i}+\lambda_{c}\cdot\Delta c_{i}$
@@
III-D Flow-based Depth Regularization
In addition to attain high reconstruction quality using matching priors and rendering depth, we also introduce the depth estimated through optical flow to calibrate the geometric structure. Since monocular depth estimation often fail to deliver accurate depth information, we incorporate epipolar priors as constraint to get depth maps explicitly.
Epipolar Line. For a point $p_{i}$ in view $i$, the corresponding epipolar line $l_{p_{i}}$ in view $j$ is defined as:
$ax+by+c=0$ (16)
$(a,b,c)^{\top}=F^{i\rightarrow j}(x_{i},y_{i},1)^{\top},$ (17)
where $F^{i\rightarrow j}$ is the fundamental matrix from view $i$ to view $j$, and $(x_{i},y_{i})$ are the coordinates of point $p_{i}$.
Flow-estimated Depth. Given a pre-trained optical flow predictor $f(\bullet)$ [69], we compute the optical flow between views $i$ and $j$ as:

### v2

@@ Abstract
keywords:
Unposed sparse-view synthesis , 3D reconstruction
journal: Pattern Recognition
Fig 1: Novel View Synthesis Comparisons.We introduce Gesplat, an efficient framework for novel view synthesis and 3D reconstruction from sparse-view unposed inputs.Compared with other pose-free methods, our method achieves higher PSNR and reconstructs more consistent geometry with finer details.
1 Introduction
@@
Neural Radiance Fields (NeRF)[1],an implicit neural representation that has shown remarkable success in reconstructing scenes and synthesizing photorealistic novel views. To improve performance under sparse views, NeRF-based methods have attempted in the aspect of pretraining [2, 3], regularization terms [4, 5], and external priors [6, 7], achieving significantly improved results. Nonetheless, these approaches still suffer from high training and rendering costs.
@@
In sparse-view settings, limited input data results in inconsistent scene geometry, and the high interdependence among Gaussian attributes forces the optimization to make a trade-off between optimizing shape and position. Some optimization-based methods [19, 20] excel in real-time scene reconstruction, but when applied to sparse-view scenarios, they usually introduce invalid geometric priors, leading to incorrect topology and scale ambiguity. Incorporating additional geometric priors is essential to constrain scene structure and improve the accuracy of 3D point positions. Depth information is a commonly used prior due to its connection between 2D images and 3D structure. Some researches [21, 22] incorporate monocular depth priors [23] as regularization, yet such handcrafted constraints often suffer from scale inaccuracy and multi-view inconsistency, which can adversely affect final rendering quality.
In this paper, we propose Gesplat, an effective framework that incorporates appropriate geometric priors to constrain the scene structure while introducing optimization and regularization techniques to refine the scene details. Inspired by [24], we adopt a hybrid Gaussian representation combining ordinary and ray-based Gaussians leveraging matching priors. By binding Gaussians to matching rays, we enhance multi-view consistency and preserve geometric structure. Furthermore, we introduce dual optimization of position and shape to stabilize training and ensure accurate surface convergence. After obtaining a relatively accurate scene representation, we employ a Graph Neural Network (GNN) [25] to further optimize Gaussian attributes. The Gaussians, treated as vertices, are connected based on spatial neighbor relationship, and learnable offsets are applied to update their attributes. To achieve more accurate depth estimation, we utilize an explicit depth computation method within an epipolar geometry framework through optical flow, following [26]. The estimated reliable depth maps are used sequentially for regularization, further improving the rendering quality. During training, we preserve joint training from [16] to optimize the Gaussian parameters and camera poses at the same time. Novel view synthesis comparisons with other pose-free methods are shown in Fig.1, which exhibit the superior performance of our method.
@@
  • 1.
    we introduce a hybrid Gaussian representation with dual optimization of position and shape based on matching priors,
  • 2.
    we design a graph-guided optimization module to refine Gaussian attributes for detailed scene recovery,
  • 3.
    we apply flow-based depth regularization to improve the quality of rendered images during training,
  • 4.
@@
2 Related Work
2.1 3D Reconstruction
@@ 2 Related Work
As a fundamental challenge in computer vision, 3D reconstruction aims to recover the 3D geometric structure of a scene from a set of input 2D images. Traditional optimization-based methods, such as Structure from Motion (SfM) [12] and Multi-View Stereo (MVS) [27, 28], rely on rigorous geometric principles to estimate accurate camera parameters and dense point clouds. COLMAP, which integrates SfM and MVS, is one of the most popular and powerful open source tools. However, its high computational cost, sensitivity to errors in SfM-estimated camera poses and sparse point clouds, and dependence on dense viewpoint coverage limit its applicability in textureless regions and areas with low overlap.The advent of deep learning has spurred the development of learning-based reconstruction methods to overcome some limitations of traditional pipelines. As a pioneering work, MVSNet [29] constructs a 3D cost volume and uses 3D convolutional neural networks to regress depth maps. Follow-up works such as R-MVSNet [30] alleviate memory consumption through recurrent networks, and CVP-MVSNet [31] introduces a coarse-to-fine pyramid structure to improve efficiency and accuracy. Nonetheless, these approaches typically assume that camera poses are known and accurate, which is usually unreliable in real-world scenarios. More recently, DUSt3R [17] notably proposes an end-to-end model for feed-forward 3D reconstruction. Unlike traditional methods that require camera calibration or viewpoint poses, DUSt3R leverages a Transformer architecture to directly estimate 3D point clouds from unconstrained image pairs, followed by global alignment to produce a complete scene representation. Its sibling, MASt3R [32], augments DUSt3R with a second network head to generate dense local features and introduces a novel matching loss for constraint. To avoid the computational cost of iterative post-optimization required by DUSt3R, VGGT [18] introduces a feed-forward network that directly generates more continuous point clouds and accurate camera poses, capable of processing more than two images simultaneously. VGGT achieves state-of-the-art reconstruction performance compared to DUSt3R and MASt3R. Inspired by this, we adopt VGGT for dense scene initialization.
2.2 Novel View Synthesis
@@
Novel view synthesis (NVS) aims to render high-quality images from unseen viewpoints. Neural Radiance Fields (NeRF) [1], which implicitly represent scenes using multi-layer perceptions (MLPs) and synthesize images via volume rendering, has become one of the most influential NVS methods. Despite achieving high-quality reconstructions, NeRF suffers from high computational costs, leading to slow training and inference. Subsequent research has improved NeRF in various aspects, $e.g.$, quality [33, 34], efficiency [35, 36], and pose-free [13, 14, 37]. More recently, 3D Gaussian Splatting (3DGS) [8] explicitly represents scenes using anisotropic 3D Gaussians [38], enabling rapid reconstruction and real-time rendering while maintaining high visual quality. 3DGS has been extensively explored in many aspects, such as [11, 20, 39, 40], demonstrating superior performance compared with NeRF-based approaches. However, conventional NVS methods heavily rely on dense input views and accurate camera poses derived from SfM, which is often impractical to attain, leading to significant performance degradation under sparse view settings.
2.3 Sparse-View Novel View Synthesis
@@
Sparse-view novel view synthesis addresses a practical but more challenging task of synthesizing novel views from a limited set of input images, where overfitting occurs due to insufficient supervision. To mitigate these issues, various strategies have been proposed. Methods like [41, 42] apply regularization techniques to reduce floating artifacts and enhance geometric consistency. [6, 43] incorporate depth priors from SfM or monocular depth estimation to supervise scene geometry, thereby reducing geometric errors and improving rendering image quality. Generalizable NeRF models like [3, 44] enable feed-forward inference, allowing fast adaptation to new scenes and improving practicality. Despite these advancements, NeRF-based methods remain computationally intensive due to volume rendering.In contrast to NeRF-based methods, several studies combined with 3DGS attempt to eliminate reconstruction inconsistency. To better capture spatial structure, [21, 22] introduce depth information into the explicit 3DGS representation. However, inaccuracies in monocular depth estimation often degrade reconstruction quality in sparse-view scenarios. [10, 45, 46] focus on robust scene initialization and removal of incorrectly positioned Gaussians. Some methods [46, 47, 48] introduce various regularization strategies to mitigate geometric degradation and suppress rendering artifacts. Alternative scene representations, such as [24, 49], are designed to produce accurate and detailed 3D scenes with enhanced geometry. To better preserve geometric structure completeness and details, we introduce a hybrid Gaussian representation to learn scene structure, with flow-based depth regularization and graph-guided optimization incorporated to refine reconstruction results in weakly textured regions.
@@
Fig 2: Overall framework of Gesplat. Given a few input images, we first generate dense point clouds and camera poses from VGGT, extract matching priors and predict optical flow. Subsequently, we randomly initialize hybrid gaussian representation, using ray-based Gaussians to optimize the gaussian position. Graph-guided optimization and joint optimization are then applied to refine the Gaussians and camera poses. Finally, we employ flow-estimated depth and matching depth as rendering geometry regularization.
3 Method
@@
3.1 Preliminary
3.1.1 Gaussian Splatting
3DGS[8] explicitly represents a scene through anisotropic 3D Gaussians, achieving high-quality scene reconstruction and real-time rendering. Each Gaussian is parameterized by: a position vector $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, spherical harmonic (SH) coefficients, and an opacity $\alpha\in[0,1)$. The influence for a point $x$ in 3D space is defined as:
@@ 3.1.1 Gaussian Splatting
3.1.2 Dense Initialization
Although 3DGS is efficient in real-time rendering, its initialization relies on a sparse point cloud and accurate camera parameters from Structure-from-Motion (SFM) [12]. This usually leads to inadequate information in sparse-view scenarios, which causes overfitting on training views and overly smooth textures. Thanks to advances in deep learning, recent learning-based 3D reconstruction methods bypass the need for explicit keypoint extraction and matching of traditional methods in an end-to-end way. Among these, DUSt3R [17] estimates 3D point clouds directly from 2D image pairs without precise camera calibration, and employs a global alignment strategy for multi-view 3D reconstruction. However, its iterative post-optimization steps result in high computation cost. To address this, we utilize VGGT [18], a feedforward neural network that produces high-fidelity dense point clouds and accurately estimates camera poses even from sparse and low-overlap datasets. This ensures geometric continuity in the reconstructed scene and provides a more robust initialization for 3DGS.
3.2 Hybrid Gaussian Representation and Optimization
@@
3.2.1 Matching Priors
Matching priors refer to ray correspondence and ray position. 1) Ray Correspondence: A pair of matching rays $\{r_{i},r_{j}\}$ from view $i$ and view $j$ corresponds to the same 3D point. Given image $I_{i}$ and $I_{j}$, matching ray pair $\{r_{i},r_{j}\}$, pixel points $\{p_{i},p_{j}\}$, intrinsic parameters $\{K_{i},K_{j}\}$, and extrinsic parameters $\{[R_{i},t_{i}],[R_{j},t_{j}]\}$, we assume that each ray intersects the surface at point $X_{i}$ and $X_{j}$, satisfying $X_{i}=X_{j}$. Thus we can get the projection from view $i$ to view $j$ as:
@@ 3.2.1 Matching Priors
3.2.2 Ray-based Gaussian
Given an image pair $I_{i}$ and $I_{j}$ with $N$ matching ray pairs $\{r_{i}^{k},r_{j}^{k}\}_{k=1}^{N}$, there exist $N$ corresponding pairs of ray-based Gaussians $\{G_{i}^{k},G_{j}^{k}\}_{k=1}^{N}$. Each primitive has similar attributes with ordinary Gaussian but differs in position representation. The new representation of position $\mu^{\prime}$ is as follows:
@@ 3.2.2 Ray-based Gaussian
3.2.3 Position Optimization
Leveraging the binding strategy between ray-based Gaussians and their matching rays, we optimize Gaussian positions using the projection relation. For a matching ray pair $\{r_{i},r_{j}\}$ with corresponding Gaussians $\{G_{i},G_{j}\}$ at positions $\mu_{i}^{\prime}=o_{i}+z_{i}d_{i}$ and $\mu_{j}^{\prime}=o_{j}+z_{j}d_{j}$, the projection error is defined as:
@@ 3.2.3 Position Optimization
3.2.4 Rendering Geometry Optimization
In addition to position, other inaccurate attributes can also affect rendering quality. To address this, we optimize the rendering geometry. Through Eq. (4), we obtain the rendering depth maps $D_{i}$ and $D_{j}$. Then we use the depth values $\{D_{i}(p_{i}),D_{j}(p_{j})\}$ in matching pixels to yield the corresponding 3D points:
@@ 3.2.4 Rendering Geometry Optimization
3.3 Graph-guided Optimization
Considering the view inconsistency under a sparse-view setting, we introduce a Graph Neural Network (GNN) to better learn the spatial geometry structure, improving the scene reconstruction quality. We utilize a graph $\mathcal{G=(V,E)}$ to capture complex geometric features for Gaussians, where vertices $\mathcal{V}=\{v_{i}\}_{i=1}^{M}=\{z_{i},s_{i},r_{i},c_{i},\alpha_{i}\}_{i=1}^{M}$ represent the Gaussian attributes, and edges $\mathcal{E}$ describe the spatial adjacency based on K nearest neighbors (KNN). The K nearest neighbors of point $\mu_{i}$ are calculated as:
@@ 3.3 Graph-guided Optimization
where $r$ is a radius threshold. Upon generating the graph network, we formulate its output as follows:
@@
$z_{i}^{\prime}$ $=z_{i}+\lambda_{z}\cdot\Delta z_{i},$ (14)
$s_{i}^{\prime}$ $=s_{i}+\lambda_{s}\cdot\Delta s_{i},$
$r_{i}^{\prime}$ $=s_{i}+\lambda_{r}\cdot\Delta r_{i},$
$c_{i}^{\prime}$ $=c_{i}+\lambda_{c}\cdot\Delta c_{i},$
@@
3.4 Flow-based Depth Regularization
In addition to attaining high reconstruction quality using matching priors and rendering depth, we also introduce the depth estimated through optical flow to calibrate the geometric structure. Since monocular depth estimation often fails to deliver accurate depth information, we incorporate epipolar priors as a constraint to explicitly compute the depth maps.
3.4.1 Epipolar Line
For a point $p_{i}$ in view $i$, the corresponding epipolar line $l_{p_{i}}$ in view $j$ is defined as:
$ax$ $+by+c=0,$ (16)
$(a,b,c)^{\top}$ $=F^{i\rightarrow j}(x_{i},y_{i},1)^{\top},$
where $F^{i\rightarrow j}$ is the fundamental matrix from view $i$ to view $j$, and $(x_{i},y_{i})$ is the coordinate of the point $p_{i}$.
3.4.2 Flow-estimated Depth
Given a pre-trained optical flow predictor $f(\bullet)$ [52], we compute the optical flow between views $i$ and $j$ as: